# MultiMind: Enhancing Werewolf Agents with Multimodal Reasoning and Theory of Mind

### v2

@@ Abstract.
Large Language Model (LLM) agents have demonstrated impressive capabilities in social deduction games (SDGs) like Werewolf, where strategic reasoning and social deception are essential. However, current approaches remain limited to textual information, ignoring crucial multimodal cues such as facial expressions and tone of voice that humans naturally use to communicate. Moreover, existing SDG agents primarily focus on inferring other players’ identities without modeling how others perceive themselves or fellow players. To address these limitations, we use One Night Ultimate Werewolf (ONUW) as a testbed and present MultiMind, the first framework integrating multimodal information into SDG agents. MultiMind processes facial expressions and vocal tones alongside verbal content, while employing a Theory of Mind (ToM) model to represent each player’s suspicion levels toward others. By combining this ToM model with Monte Carlo Tree Search (MCTS), our agent identifies communication strategies that minimize suspicion directed at itself. Through comprehensive evaluation in both agent-versus-agent simulations and studies with human players, we demonstrate MultiMind’s superior performance in gameplay. Our work presents a significant advancement toward LLM agents capable of human-like social reasoning across multimodal domains.
@@
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06ccs: Computing methodologies Natural language generationccs: Computing methodologies Activity recognition and understandingccs: Computing methodologies Speech recognition
@@ 1. Introduction
Large Language Model (LLM) agents have made significant progress in simulating human-like decision-making and social behavior across various domains. Recent research has demonstrated their ability to play games such as StarCraft (Ma et al., 2024; Shao et al., 2024) and MineCraft (Wang et al., 2024; Qin et al., 2024; Li et al., 2024). Within this broad research landscape, a promising direction has developed focusing on games that require not just strategic thinking but also sophisticated social reasoning. Social deduction games (SDGs) such as Werewolf (Xu et al., 2023, 2024; Jin et al., 2024; Wu et al., 2024b), Avalon (Light et al., 2023; Wang et al., 2023; Shi et al., 2023; Lan et al., 2024), and Jubensha (Liang et al., 2025; Wu et al., 2024a) have proven to be challenging testbeds for LLM agent research. These games demand nuanced reasoning and strategic communication based on hidden information, which are skills that more closely resemble human social intelligence.
@@ 4.3. Ablation Study
Figure 6. A dialogue between a MultiMind agent and a human participant. The left panel shows the agent’s ToM assessments and decision-making process. The right panel displays the dialogue from the human, including facial and speech emotion labels detected by the Perceiver.
@@ 4.5. Multimodal Cues in ToM Modeling
We also provide a case study in Figure 6 that illustrates how our agent integrates multimodal cues when interacting with human participants. In this case, MultiMind successfully misleads the human player, ultimately leading to a vote against Player 2. The integration of the multimodal signals enables more nuanced understanding of human intentions and mental states, allowing the agent to adapt its strategy accordingly in the social deduction game context.

### v4

@@ Abstract.
Large Language Model (LLM) agents have demonstrated impressive capabilities in social deduction games (SDGs) like Werewolf, where strategic reasoning and social deception are essential. However, current approaches remain limited to textual information, ignoring crucial multimodal cues such as facial expressions and tone of voice that humans naturally use to communicate. Moreover, existing SDG agents primarily focus on inferring other players’ identities without modeling how others perceive themselves or fellow players. To address these limitations, we use One Night Ultimate Werewolf (ONUW) as a testbed and present MultiMind, the first framework integrating multimodal information into SDG agents. MultiMind processes facial expressions and vocal tones alongside verbal content, while employing a Theory of Mind (ToM) model to represent each player’s suspicion levels toward others. By combining this ToM model with Monte Carlo Tree Search (MCTS), our agent identifies communication strategies that minimize suspicion directed at itself. Through comprehensive evaluation in both agent-versus-agent simulations and studies with human players, we demonstrate MultiMind’s superior performance in gameplay. Our work presents a significant advancement toward LLM agents capable of human-like social reasoning across multimodal domains. Our code is available at https://github.com/CjangCjengh/onuw.
@@
journalyear: 2025copyright: acmlicenseddoi: 10.1145/3746027.3755752conference: Proceedings of the 33rd ACM International Conference on Multimedia; October 27–31, 2025; Dublin, Ireland.booktitle: Proceedings of the 33rd ACM International Conference on Multimedia ($M$ ’25), October 27–31, 2025, Dublin, Irelandisbn: 979-8-4007-2035-2/2025/10ccs: Computing methodologies Natural language generationccs: Computing methodologies Activity recognition and understandingccs: Computing methodologies Speech recognition
@@ 1. Introduction
Large Language Model (LLM) agents have made significant progress in simulating human-like decision-making and social behavior across various domains. Recent research has demonstrated their ability to play games such as StarCraft (Ma et al., 2024; Shao et al., 2024) and MineCraft (Wang et al., 2024; Qin et al., 2024; Li et al., 2024). Within this broad research landscape, a promising direction has developed focusing on games that require not just strategic thinking but also sophisticated social reasoning. Social deduction games (SDGs) such as Werewolf (Zhang et al., 2025; Xu et al., 2023, 2024; Wu et al., 2024b), Avalon (Light et al., 2023; Wang et al., 2023; Shi et al., 2023; Lan et al., 2024), and Jubensha (Liang et al., 2025; Wu et al., 2024a) have proven to be challenging testbeds for LLM agent research. These games demand nuanced reasoning and strategic communication based on hidden information, which are skills that more closely resemble human social intelligence.
@@ 5. Conclusion
Acknowledgements.
This research is supported by Tencent Rhino-Bird Focused Research Program, National Natural Science Foundation of China (No. 62406267), and the Guangzhou Municipal Science and Technology Project (No. 2025A04J4070).